OpenStudy (anonymous):

solve using the substitution method 2x+5y=14 -5x+y=19

OpenStudy (anonymous):

You posted this before. I answered it there. Do you have any questions as to how I did it?

OpenStudy (anonymous):

I didn't understand at all and I should have 2 numbers for ans that look like this. ( x,y) but with numbers obviously

OpenStudy (anonymous):

I came up with solutions: x=-3 and y=4 You can write them (-3,4)

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Step 1: Pick one equation out of the two and solve for one of the variables I picked the second equation and I picked the letter y the second equation -5x+y=19 Moving the x to the other side of the equal sign, you get: y=19+5x Now you know what y is, but that's not quite a solution yet. So now this value of y can be substituted in the other equation Since this value came from the second equation, we take the first equation: 2x+5y=14 Only instead of the letter y, we will use the equation we got in step 1 (y=19+5x) So we can rewrite equation 1 with the new value of y 2x+5(19+5x)=14 We multiply out the 5*(19+5x) 2x+95+25x=14 We collect the x and move the 95 to the other side 27x=-81 Now just divide both sides by 27 x=-3 You have one of the solutions!! Now just substitute the value of x in either equation: (Here I'll pick the first one) 2x+5y=14 2(-3)+5y=14 -6+5y=14 MOving the 6 to the other side: 5y=14+6 or 5y=20 And finally dividing by 5 on both sides y=4 Now you have both solutions x=-3 and y=4 or (-3,4) Good luck.